An electric field of 9.00×105 V/m is desired between two parallel plates, each of area 45.0 cm2 and separated by 2.45 mm of air. What charge must be on each plate?
Answer #1

Capacitance, keo A C =-

For air, k=1

So c_1*8.854 * 10-14 * 0.0045m 0.00245m

8.854*10-12 + 0.0045 0.00245

C = 1.626 * 10-11 F

=======================

Electric Field, E = = 9.00 * 10%V/m

” = 9.00 * 10%v/m

\frac{V}{0.00245m}=9.00*10^{5}V/m

V = 9.00 * 10% * 0.00245

{\color{Red} V=2205V}

==============

Charge, Q=CV = 1.626 * 10-11 F * 22051

ANSWER: Q = 3.58533 * 10-80

================================